The rate constant of a reaction is 4.75 × 10−5 L/mol·s at 195 °C and 9.85 × 10−3 L/mol·s at 258 °C. What is the activation energy of the reaction? Enter your answer in scientific notation.
Answer #1

Answer

175.1kJ/mol

Explanation

From Arrhenius equation we get the following relation

log(k1/k2) = Ea/2.303R(1/T2 - 1/T1)

where,

k1 = rate constant 1, 4.75×10-5M-1s-1

k2 = rate constant 2 , 9.85 ×10-3M-1s-1

T1 = temperature 1, 195℃ = 468.15K

T2 = temperature 2 , 258℃ = 531.15K

R = gas constant , 8.314J/K mol

Ea = activation energy

substituting the values

log(4.75×10-5M-1s-1/9.85×10-3M-1s-1) = (Ea/2.303×8.314J/Kmol)(1/531.15K - 1/468.15K)

- 2.3168= (Ea/19.147J/molK) (-0.0002534K-1)

Ea/19.147J/mol = 9142.9

Ea = 175.1kJ/mol